# Internal Dynamics of Multiple Populations in 28 Galactic Globular Clusters: A Wide-Field study with Gaia and the Hubble Space Telescope

### v1

@@ 4 Discussion
To investigate how these dynamical differences depend on the clusters’ properties evolutionary state, we divided our sample of 28 clusters into different groups based on various internal and external factors: dynamical ages (determined as the ratio between cluster age and half-mass relaxation time, see e.g. Libralato et al., 2022), tidal filling factor defined as (Baumgardt et al., 2010; Shin et al., 2013) with being the Jacobi radius derived in Balbinot & Gieles (2018)999Clusters with are considered in “isolated regime” or tidally underfilling clusters, while clusters with are in “tidal regime”, or tidally filling., spatial mixing of MPs (as determined in Leitinger et al., 2023, and Jang et al. in preparation), tidal interaction with the Milky Way (determined on the basis of the peri-Galactic radius, see e.g. Milone et al., 2020; Libralato et al., 2023)101010Consistently with literature works, we divide clusters with , clusters’ origin (e.g. in situ or accreted) as identified in Massari et al. (2019) and escape velocity (, see e.g. Baumgardt & Hilker, 2018; Mastrobuono-Battisti & Perets, 2021)111111See e.g. the analysis of Baumgardt & Hilker (2018) and Mastrobuono-Battisti & Perets (2021) and the connection between the cluster’s and MPs. The value of the limiting escape velocity, i.e. , is determined examining the distribution of escape velocities from Baumgardt & Hilker (2018).. The groups are shown in the table provided in App. C. We find worth mentioning that different groups have overlap, as some clusters’ properties are not independent one from the other (e.g. the spatial mixing of 1P and 2P is connected with the cluster dynamical age Dalessandro et al., 2019; Leitinger et al., 2023). Additionally, we also divided clusters based on their internal bulk rotation (as determined in Vasiliev & Baumgardt, 2021), but, to keep the figure simpler, we not show the results in Fig. 4. Nonetheless, the difference between 1P and 2P anisotropy profiles for the groups and the statistical significance are included in App. C.
@@
We find significant differences, i.e. above the 3 level, in the average anisotropy profiles of many of the analyzed groups. Overall, we find that the 2P is more radially anisotropic beyond , while 1P is isotropic and become slightly tangentially anisotropic beyond 2-3. There are important differences in the 1P-2P relative dynamics among some of the analyzed groups. For example, while dynamically young clusters and those with centrally concentrated 2P stars show clearly different dispersion and anisotropy profiles, these differences are less pronounced in intermediate-age and mixed-population clusters. These findings are qualitatively consistent with the simulations by Tiongco et al. (2019, see, e.g., their Fig. 9) and Vesperini et al. (2021, see, e.g., their Figs. 13-17), which suggest that due to different initial spatial configurations (such as centrally concentrated 2P stars) or initial isotropy, 1P stars tend to develop isotropic motion, while 2P stars display more radially anisotropic motion, especially in the outer regions. Over time, these dynamical differences gradually diminish as spatial and dynamical mixing of MPs occurs.
The comparison between clusters with small and large peri-Galactic radii (shown in the first two columns of the bottom row of Fig. 4) reveals some interesting patterns. In clusters with small peri-Galactic radii, 1P stars exhibit tangential anisotropy beyond , while in clusters with larger peri-Galactic radii, 1P stars are either isotropic or slightly radially anisotropic. In both cases, 2P stars become increasingly radially anisotropic as they move outward. These findings suggest that the interaction with the Milky Way’s tidal field plays a significant role in the dynamical evolution of MPs. Moreover, in clusters with small peri-Galactic radii, 1P stars appear to behave like tidally filling systems, where the outer regions develop slight tangential anisotropy, whereas 2P stars resemble tidally underfilling systems with radial anisotropy. In contrast, weaker interactions with the Milky Way’s tidal field result in a more isotropic 1P population. This scenario aligns qualitatively with the theoretical models presented in Tiongco et al. (2019); Hypki et al. (2022); Pavlík et al. (2024).
@@
In summary, the results presented in Fig. 4 indicate that clusters in a less dynamically evolved state show significant dynamical differences among MPs. Additionally, the interaction with the Galaxy appears to play a crucial role in shaping the evolution of different populations. These findings align with the conclusions of Libralato et al. (2023) regarding the internal dynamics of MPs in the innermost regions.
@@ 5 Summary and conclusion
In this study, we present the first comprehensive investigation of the internal dynamics of multiple populations across a wide field of view, from the innermost arcmin to the clusters’ outskirts, in a large sample of 28 Galactic GCs. Using HST, ground-based and XP synthethic photometry, we identified first- and second-population stars among RGB stars from the innermost regions to the clusters’ outskirts. To achieve this, we exploited the pseudo CMD dubbed Chromosome maps as well as the index, where 1P and 2P stars form well-separated groups. The internal dynamics of MPs was investigated using HST and Gaia DR3 proper motions, allowing us to determine the mean, dispersion, and anisotropy profiles as function of distance from the cluster center. The analyzed cluster regions range from 0.2 to more than 10 . The analysis presented in this work confirms and extends the results of previous studies based on HST and Gaia proper motions (see e.g. Richer et al., 2013; Bellini et al., 2015, 2018; Milone et al., 2018; Cordoni et al., 2020a, b; Libralato et al., 2023; Cadelano et al., 2024).
@@
Studying the global trends in dispersion and anisotropy profiles, derived by combining the dynamical profiles of various clusters, we observe significant differences between 1P and 2P stars, especially outside where the differences are significant beyond the 3 level. In dynamically young clusters, 1P stars are isotropic in the inner regions but become slightly tangentially anisotropic toward the outskirts. In contrast, 2P stars are isotropic in the cluster centers and become radially anisotropic in the outer regions. These patterns are also evident in clusters where 2P stars are more centrally concentrated and in clusters with escape velocities exceeding . However, dynamical evolved clusters, with spatially mixed MPs and lower escape velocities do not exhibit these dynamical differences. Rotating/non-rotating clusters, and in situ/accreted clusters show similar relative differences between 1P and 2P stars.
We also explore the influence of the Milky Way’s tidal field on the dynamical properties of MPs by analyzing clusters with different peri-Galactic radii. Our findings reveal a strong connection between the dynamical behavior of 1P stars and the strength of the Milky Way’s tidal field. In clusters with orbits closer to the Galactic center, where the tidal field is stronger, 1P stars tend to exhibit tangential anisotropy beyond 1-2. Conversely, in clusters with weaker interactions with the Milky Way’s tidal field, 1P stars display isotropic motion. This suggests that the Milky Way’s tidal field plays a crucial role in the dynamical evolution of MPs. Further supporting this conclusion, our analysis shows that tidally underfilling and filling clusters exhibit distinct relative patterns in the dynamical profiles of 1P and 2P stars.
@@ Acknowledgments
SJ acknowledges support from the NRF of Korea (2022R1A2C3002992, 2022R1A6A1A03053472). EPL acknowledges support from the “Science & Technology Champion Project” (202005AB160002) and from the “Top Team Project” (202305AT350002), all funded by the “Yunnan Revitalization Talent Support Program”. TZ acknowledges funding from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie Grant Agreement No. 101034319 and from the European Union – Next Generation EU. This work has received funding from “PRIN 2022 2022MMEB9W - Understanding the formation of globular clusters with their multiple stellar generations” (PI Anna F. Marino), and from INAF Research GTO-Grant Normal RSN2-1.05.12.05.10 - (ref. Anna F. Marino) of the “Bando INAF per il Finanziamento della Ricerca Fondamentale 2022”. EL acknowledges financial support from the European Research Council for the ERC Consolidator grant DEMOBLACK, under contract no. 770017
@@
Relevant data underlying this work is available in the article. All other data will be shared upon reasonable request to the corresponding author.
@@ Appendix A Selection of Multiple Populations
Concerning ground-based observations, we first used the selection in the ChM whenever available, and we adopted the selection for stars without ChM information. MPs outside the orange circles are selected by means of Gaia XP synthethic photometry. We remind here that Gaia XP synthethic photometry is only available for four clusters, namely NGC 0104, NGC 3201, NGC 6121 and NGC 6752 (see Mehta et al., 2024, for a complete discussion).
@@ Appendix B Individual profiles of all clusters
Cluster | Dynamical age | MPs | Rotation | Peri radius | Origin | Filling factor | |
NGC 0104 | young | p2conc | rot | outer | high | in-situ | underfilling |
NGC 0288 | young | mixed | non rot | inner | low | accreted | filling |
NGC 1261 | young | mixed | non rot | inner | high | accreted | underfilling |
NGC 1851 | young | p2conc | rot | inner | high | accreted | underfilling |
NGC 1904 | unknown | p2conc | rot | inner | unknown | accreted | unknown |
NGC 2298 | intermediate-old | p2conc | non rot | inner | low | accreted | underfilling |
NGC 2808 | young | p2conc | rot | inner | high | accreted | underfilling |
NGC 3201 | young | p2conc | non rot | outer | low | accreted | underfilling |
NGC 4147 | unknown | mixed | non rot | inner | unknown | accreted | unknown |
NGC 4590 | young | mixed | non rot | outer | low | accreted | underfilling |
NGC 4833 | young | p2conc | non rot | inner | low | accreted | filling |
NGC 5053 | young | p1conc | non rot | outer | low | accreted | filling |
NGC 5904 | young | p2conc | rot | inner | high | accreted | underfilling |
NGC 5272 | young | p2conc | rot | outer | high | accreted | underfilling |
NGC 5927 | young | mixed | non rot | outer | high | in-situ | filling |
NGC 6121 | intermediate-old | mixed | non rot | inner | low | in-situ | filling |
NGC 6205 | young | mixed | non rot | inner | high | accreted | underfilling |
NGC 6218 | intermediate-old | mixed | rot | inner | low | in-situ | filling |
NGC 6254 | intermediate-old | p2conc | non rot | inner | high | in-situ | filling |
NGC 6366 | intermediate-old | mixed | non rot | inner | low | in-situ | filling |
NGC 6712 | unknown | p1conc | non rot | inner | unknown | in-situ | unknown |
NGC 6752 | intermediate-old | mixed | rot | inner | high | in-situ | underfilling |
NGC 6809 | young | mixed | norot | inner | low | in-situ | filling |
NGC 6838 | intermediate-old | mixed | non rot | outer | low | in-situ | filling |
NGC 6934 | young | mixed | norot | inner | high | in-situ | underfilling |
NGC 6981 | intermediate-old | p2conc | non rot | inner | low | accreted | underfilling |
NGC 7492 | unknown | mixed | non rot | inner | unknown | accreted | unknown |
NGC 7089 | young | mixed | rot | inner | high | accreted | underfilling |

### v2

@@ 4 Discussion
To investigate how these dynamical differences depend on the clusters’ properties evolutionary state, we divided our sample of 28 clusters into different groups based on various internal and external factors: dynamical ages (determined as the ratio between cluster age and half-mass relaxation time, see e.g. Libralato et al., 2022), tidal filling factor defined as (Baumgardt et al., 2010; Shin et al., 2013) with being the Jacobi radius derived in Balbinot & Gieles (2018)999Clusters with are considered in “isolated regime” or tidally underfilling, while clusters with are in “tidal regime”, or tidally filling., spatial mixing of MPs (as determined in Leitinger et al., 2023, and Jang et al., in review), tidal interaction with the Milky Way (determined on the basis of the peri-Galactic radius, see e.g. Milone et al., 2020; Libralato et al., 2023)101010Consistently with literature works, we divide clusters with , clusters’ origin (e.g. in situ or accreted) as identified in Massari et al. (2019) and escape velocity (, see e.g. Baumgardt & Hilker, 2018; Mastrobuono-Battisti & Perets, 2021)111111See e.g. the analysis of Baumgardt & Hilker (2018) and Mastrobuono-Battisti & Perets (2021) and the connection between the cluster’s and MPs. The value of the limiting escape velocity, i.e. , is determined examining the distribution of escape velocities from Baumgardt & Hilker (2018).. The groups are shown in the table provided in App. C. We find worth mentioning that different groups have overlap, as some clusters’ properties are not independent one from the other (e.g. the spatial mixing of 1P and 2P is connected with the cluster dynamical age Dalessandro et al., 2019; Leitinger et al., 2023). Additionally, we also divided clusters based on their internal rotation (as determined in Vasiliev & Baumgardt, 2021), but, to keep the figure simpler, we do not show the results in Fig. 4. Nonetheless, the difference between 1P and 2P anisotropy profiles for the groups, along with the statistical significance, is included in Fig. 13.
@@
We find significant differences, i.e. above the 3 level, in the average anisotropy profiles of many of the analyzed groups. Overall, we find that the 2P is more radially anisotropic beyond , while 1P is isotropic and become slightly tangentially anisotropic beyond 2-3. There are important differences in the 1P-2P relative dynamics among some of the analyzed groups. For example, while dynamically young clusters and those with centrally concentrated 2P stars show clearly different dispersion and anisotropy profiles, these differences are less pronounced in intermediate-age and mixed-population clusters. Indeed, dynamical profile of 1P and 2P in dynamically old clusters are consistent within . These findings are qualitatively consistent with the simulations by Tiongco et al. (2019, see, e.g., their Fig. 9) and Vesperini et al. (2021, see, e.g., their Figs. 13-17), which suggest that due to different initial spatial configurations (such as centrally concentrated 2P stars) or initial isotropy, 1P stars tend to develop isotropic motion, while 2P stars display more radially anisotropic motion, especially in the outer regions. Over time, these dynamical differences gradually diminish as spatial and dynamical mixing of MPs occurs.
The comparison between clusters with small and large peri-Galactic radii (referred to as “inner” and “outer” clusters, respectively) reveals distinct dynamical patterns (Fig. 4, bottom row). In inner clusters, 1P stars exhibit tangential anisotropy beyond , while in outer clusters, 1P is either isotropic or slightly radially anisotropic between 1 and 2 . In both cases, 2P stars become increasingly radially anisotropic as they move outward. Although the differences in 1P dynamics between inner and outer clusters are only marginally significant in 1-2 region (ranging between 1–2), these findings suggest that the Milky Way’s tidal field significantly influences the dynamical evolution of MPs. Moreover, in clusters with small peri-Galactic radii, 1P stars appear to behave like tidally filling systems, where the outer regions develop slight tangential anisotropy, whereas 2P stars resemble tidally underfilling systems with radial anisotropy. In contrast, weaker interactions with the Milky Way’s tidal field result in a more isotropic 1P population. This scenario aligns qualitatively with the theoretical models presented in Tiongco et al. (2019); Hypki et al. (2022); Pavlík et al. (2024).
@@
In summary, the results presented in Fig. 4 indicate that clusters in a less dynamically evolved state show significant dynamical differences among MPs. Additionally, the interaction with the Galaxy appears to play a crucial role in shaping the evolution of different populations. These findings align with the conclusions of Libralato et al. (2023) regarding the internal dynamics of MPs in the innermost regions. However, while our results clearly support the presence of a common dynamical pattern, it is important to note that a small number of clusters deviate from this trend, exhibiting distinct dynamical properties. The nature of these differences remains unclear and may be influenced by specific cluster properties and their evolutionary histories.
4.1 Comparison with the literature
In recent years, numerous studies have investigated the dynamics of MPs in GCs (see references in Sec. 1). Here, we provide a comparison with the findings of Libralato et al. (2023); Dalessandro et al. (2024); Leitinger et al. (2024), who analyzed 56, 16 and 30 GCs, respectively. Our conclusions regarding the dynamical properties of 1P/2P stars in both dynamically young and old clusters are consistent with the results of Libralato et al. (2023) and Dalessandro et al. (2024). Specifically, our findings support the presence of a general trend in the dynamical properties of 1P and 2P stars, which suggests a common underlying dynamical pattern. On a cluster-by-cluster basis, our anisotropy profiles qualitatively align with the parameters reported by Dalessandro et al., with the exception of NGC 5927. However, direct comparisons are limited, as their study employs analytical fits to model the dispersion and anisotropy profiles, whereas we utilize profiles derived from radial bins. Similarly, we cannot directly compare results on the internal rotation of individual clusters or the global rotation profile presented by Dalessandro et al. (2024), as we chose not to fit the tangential profiles with analytical models. A direct comparison with Leitinger et al. (2024) is also challenging, as they focus more on MPs rotation and present anisotropy profiles for all cluster stars without separating MPs.
@@ 5 Summary and conclusion
In this study, we present a detailed investigation of the internal dynamics of multiple populations across a wide field of view, from the innermost arcmin to the clusters’ outskirts, in a large sample of 28 Galactic GCs. Using HST, ground-based and XP synthethic photometry, we identified first- and second-population stars. To achieve this, we exploited the pseudo CMD dubbed Chromosome maps as well as the index, where 1P and 2P stars form well-separated groups. The internal dynamics of MPs was investigated using HST and Gaia DR3 proper motions, allowing us to determine the mean, dispersion, and anisotropy profiles as function of distance from the cluster center. The analyzed cluster regions range from 0.2 to more than 10 . The analysis presented in this work confirms and extends the results of previous studies based on HST and Gaia proper motions (see e.g. Richer et al., 2013; Bellini et al., 2015, 2018; Milone et al., 2018; Cordoni et al., 2020a, b; Libralato et al., 2023; Cadelano et al., 2024; Dalessandro et al., 2024).
@@
Studying the global trends in dispersion and anisotropy profiles, derived by combining individual clusters’ catalogs (see Sec. 3.2 for a description of the normalization), we observe significant differences between 1P and 2P stars, especially outside where the differences are significant beyond the 3 level.
In dynamically young clusters, 1P stars are isotropic in the inner regions but become slightly tangentially anisotropic toward the outskirts. In contrast, 2P stars are isotropic in the cluster centers and become radially anisotropic in the outer regions. This result is in agreement with the results presented in Dalessandro et al. (2024) for a sample of 16 GCs. These patterns are also evident in clusters where 2P stars are more centrally concentrated and in clusters with escape velocities exceeding . However, dynamical evolved clusters, with spatially mixed MPs and lower escape velocities do not exhibit these dynamical differences. Rotating/non-rotating clusters, and in situ/accreted clusters show similar relative differences between 1P and 2P stars.
We also explore the influence of the Milky Way’s tidal field on the dynamical properties of MPs by analyzing clusters with different peri-Galactic radii. Our findings reveal a possible connection between the dynamical behavior of 1P stars and the strength of the Milky Way’s tidal field. In clusters with orbits closer to the Galactic center, where the tidal field is stronger, 1P stars tend to exhibit tangential anisotropy beyond 1-2. Conversely, clusters with weaker interactions with the Milky Way’s tidal field, 1P stars display isotropic motion. This suggests that the Milky Way’s tidal field plays an important role in the dynamical evolution of MPs. Further supporting this conclusion, our analysis shows that tidally underfilling and filling clusters exhibit distinct relative patterns in the dynamical profiles of 1P and 2P stars.
@@ Acknowledgments
SJ acknowledges support from the NRF of Korea (2022R1A2C3002992, 2022R1A6A1A03053472). EPL acknowledges support from the “Science & Technology Champion Project” (202005AB160002) and from the “Top Team Project” (202305AT350002), all funded by the “Yunnan Revitalization Talent Support Program”. TZ acknowledges funding from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie Grant Agreement No. 101034319 and from the European Union – Next Generation EU. This work has received funding from “PRIN 2022 2022MMEB9W - Understanding the formation of globular clusters with their multiple stellar generations” (PI Anna F. Marino), and from INAF Research GTO-Grant Normal RSN2-1.05.12.05.10 - (ref. Anna F. Marino) of the “Bando INAF per il Finanziamento della Ricerca Fondamentale 2022”. EL acknowledges financial support from the European Research Council for the ERC Consolidator grant DEMOBLACK, under contract no. 770017. AMB acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 895174.
@@
Relevant data underlying this work is available in the article. Additional figures and tests mentioned in the artcile are available on GitHub and as a Streamlit application. All other data will be shared upon reasonable request to the corresponding author.
@@ Appendix A Selection of Multiple Populations
Concerning ground-based observations, we first used the selection in the ChM whenever available, and we adopted the selection for stars without ChM information. MPs outside the orange circles are selected by means of Gaia XP synthethic photometry. We remind here that Gaia XP synthethic photometry is only available for four clusters, namely NGC 0104, NGC 3201, NGC 6121 and NGC 6752 (see Mehta et al., 2024, for a complete discussion). For Type II clusters in our sample, we show anomalous stars as orange points in HST and ground-based ChMs.
@@ Appendix B Individual profiles of all clusters
Cluster | Dynamical age | MPs | Rotation | Peri radius | Origin | Filling factor | ||||
NGC 0104 | young | p2conc | rot | outer | high | in-situ | underfilling | 314/827 | 1564/1252 | –/– |
NGC 0288 | young | mixed | non rot | inner | low | accreted | filling | 118/146 | 100/135 | –/– |
NGC 1261 | young | mixed | non rot | inner | high | accreted | underfilling | 344/62 | 508/82 | 33/14 |
NGC 1851 | young | p2conc | rot | inner | high | accreted | underfilling | 278/85 | 572/103 | 284/71 |
NGC 1904 | unknown | p2conc | rot | inner | unknown | accreted | unknown | –/56 | –/128 | –/– |
NGC 2298 | intermediate-old | p2conc | non rot | inner | low | accreted | underfilling | 59/26 | 87/31 | –/– |
NGC 2808 | young | p2conc | rot | inner | high | accreted | underfilling | 449/351 | 1365/397 | –/– |
NGC 3201 | young | p2conc | non rot | outer | low | accreted | underfilling | 66/352 | 96/284 | –/– |
NGC 4147 | unknown | mixed | non rot | inner | unknown | accreted | unknown | –/19 | –/37 | –/– |
NGC 4590 | young | mixed | non rot | outer | low | accreted | underfilling | 106/23 | 219/60 | –/– |
NGC 4833 | young | p2conc | non rot | inner | low | accreted | filling | 156/45 | 225/62 | –/– |
NGC 5053 | young | p1conc | non rot | outer | low | accreted | filling | 32/14 | 23/60 | –/– |
NGC 5904 | young | p2conc | rot | inner | high | accreted | underfilling | 364/306 | 871/318 | –/– |
NGC 5272 | young | p2conc | rot | outer | high | accreted | underfilling | 206/298 | 751/475 | –/– |
NGC 5927 | young | mixed | non rot | outer | high | in-situ | filling | 141/47 | 441/87 | –/– |
NGC 6121 | intermediate-old | mixed | non rot | inner | low | in-situ | filling | 38/104 | 94/240 | –/– |
NGC 6205 | young | mixed | non rot | inner | high | accreted | underfilling | 221/230 | 1009/379 | –/– |
NGC 6218 | intermediate-old | mixed | rot | inner | low | in-situ | filling | 131/126 | 179/202 | –/– |
NGC 6254 | intermediate-old | p2conc | non rot | inner | high | in-situ | filling | –/218 | –/255 | –/– |
NGC 6366 | intermediate-old | mixed | non rot | inner | low | in-situ | filling | 29/55 | 43/54 | –/– |
NGC 6712 | unknown | p1conc | non rot | inner | unknown | in-situ | unknown | –/49 | –/42 | –/– |
NGC 6752 | intermediate-old | mixed | rot | inner | high | in-situ | underfilling | 103/178 | 265/360 | –/– |
NGC 6809 | young | mixed | norot | inner | low | in-situ | filling | 47/60 | 117/118 | –/– |
NGC 6838 | intermediate-old | mixed | non rot | outer | low | in-situ | filling | 67/78 | 32/50 | –/– |
NGC 6934 | young | mixed | norot | inner | high | in-situ | underfilling | 205/25 | 364/89 | 30/– |
NGC 6981 | intermediate-old | p2conc | non rot | inner | low | accreted | underfilling | 201/53 | 186/50 | –/– |
NGC 7089 | young | mixed | rot | inner | high | accreted | underfilling | 273/165 | 941/329 | 57/20 |
NGC 7492 | unknown | mixed | non rot | inner | unknown | accreted | unknown | –/27 | –/49 | –/– |